The wheels of a car have radius 15 in, and are rotating at 800 rpm. Find the speed of the car in mi/h
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

C=2PIR
C=2*3.14*15
C=6.28*15
C=94.2 INCHES PER ROTATION.
94.2*800=75,360 IN/MIN.
75,360/12=6,280 FT/MIN.
6,280*60=376,800 FT/HR.
376,800/5280=71.36 MPH ANS.

Answer by Alan3354(69443)

The wheels of a car have radius 15 in, and are rotating at 800 rpm. Find the speed of the car in mi/h
-------------------
The car moves 1 circumference per revolution
C = 2*15*pi = 30pi inches per rev
@ 800 rpm, that's 24000pi inches/minute
= 2000pi ft/minute
= 120,000pi ft/hour
= 120,000pi/5280 mph
= 71.4 mph